OpenStudy (karon86):

3x-y=12, What is the value of 8^x/2^y?

OpenStudy (karon86):

\[\frac{ 8^x }{ 2^y }\]

OpenStudy (karon86):

What is the value?

OpenStudy (karon86):

A. 2^12 B. 4^4 C. 8^2 D. Value cannot be found

OpenStudy (karon86):

YAY :D

OpenStudy (anonymous):

i do not know @KendrickLamar2014

OpenStudy (karon86):

XDDD

OpenStudy (karon86):

It's okay. I don't either XD

OpenStudy (karon86):

I substituted y for (-12+3x)

OpenStudy (karon86):

And I have 3x-(-12x+3x)=12 But if I was to continue to solve I wouldn't get the answer

OpenStudy (karon86):

I meant y for (-12+3x)

OpenStudy (anonymous):

is this 8^x/2^y 8 times x divided by 2 times y?

OpenStudy (karon86):

yes

OpenStudy (anonymous):

ok then

OpenStudy (karon86):

But we have to solve for x and y :I

OpenStudy (anonymous):

im gonna ask one of my friends for help one sec

OpenStudy (karon86):

okay. I'll be waiting ^_^

OpenStudy (anonymous):

3x-y=12, What is the value of 8^x/2^y? he said 4xy

OpenStudy (karon86):

how did he get that?

OpenStudy (karon86):

I got \[4 ^{1/4}\]

OpenStudy (anonymous):

I gtg tty guys later

OpenStudy (karon86):

my answer is no where near the answer though XD

OpenStudy (karon86):

Hi :D

OpenStudy (karon86):

@Luigi0210

OpenStudy (karon86):

Hi :D

OpenStudy (karon86):

Don't bother many have tried :(

OpenStudy (karon86):

Okay no pleaaaseee help me!

OpenStudy (ikram002p):

Hint :- 3x-y=12, take both side to the power 2 \(\Large 2^{3x-y}=2^{12}\)

OpenStudy (karon86):

O_O

OpenStudy (karon86):

What?

OpenStudy (karon86):

I am giving you a medal if you got this right okay let me see

OpenStudy (karon86):

Um. I don't get it

OpenStudy (ikram002p):

Oh dear... i don't care for medals just trying to help

OpenStudy (karon86):

Oh thank you anyway. Um. I don't get it

OpenStudy (ikram002p):

here again